OpenStudy (nannie):

Show me how to solve this: 8/v=9/v-1/3

OpenStudy (nikita2):

8/v=9/v-1/3 | *v multiply 8 = 9 - 1/3v 1/3v = 9 - 8 1/3v = 1 | *3 v = 3

hero (hero):

Nikita's pic looks like Andy Rodd ick mixed with Andy Murray